Solve each system by addition, if possible
Having trouble understanding this.
x + y = 1
x - y = 7
Do I add it together like this:
x + y = 1
x - y = 7
--------------
2x = 8
Then, should I divide it like this:
8/2x = 4

Found 3 solutions by solver91311, richwmiller, ankor@dixie-net.com:
Answer by solver91311(24713)



You have the correct process and got to half of the right answer, but you expressed the last step incorrectly.





But that is only half of the problem. What is the value of ?

Now that you have found the value of , substitute back into either of your original equations:







Finally, check the coordinates of your ordered pair solution set in the other equation:





Checks.

Now you can report that the solution set of the given system contains a unique ordered pair, and the solution set is

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by richwmiller(17219)

close
2x=8
x=8/2
x=4
then go back and find y
x+y=1
4+y=1
y=-3

Answer by ankor@dixie-net.com(22740)

Solve each system by addition, if possible
Having trouble understanding this.
x + y = 1
x - y = 7
Do I add it together like this:
x + y = 1
x - y = 7
--------------
2x = 8
Then, This
x = 8/2
x = 4
Find y using the 1st equation
4 + y = 1
y = 1 - 4
y = -3
Check the solutions in the 2nd equation
4 - (-3) =
Minus a minus is plus, therefore
4 + 3 = 7